# LearnLM: Improving Gemini for Learning

### v2

@@ 4.1 Safety evaluation
Similar to the process described in our initial tech report [1] and the Gemini tech reports [3, 17], safety, responsibility, and assurance evaluations were carried out on LearnLM in collaboration with Google DeepMind’s Responsible Development and Innovation team and Google’s Trust and Safety team to ensure adherence to Gemini’s model policy as well as a learning-specific model policy.
@@
Due to our reframing in terms of pedagogical instruction following and our decision to co-train (see Section 2), our training and safety evaluation procedure is now fully aligned with Gemini 1.5. See Table 45 in Appendix 12 of the Gemini 1.5 report [3] for a model card. For details on learning-specific dataset curation and safety evaluations, and a discussion of ethical risks and limitations, see Section 2 and the original LearnLM tech report [1], including the model card presented in Appendix A therein.
@@ 5 Conclusion
We have described our motivation and approach to improving foundation models for learning use cases, which relies on System Instructions to condition desired behavior. We updated Gemini’s post-training mixture to add demonstration data (via SFT) and human preference data (via a Reward Model and RLHF) to teach the model to follow a range of pedagogical instructions. We then evaluated the resulting LearnLM model alongside comparable models, showing significant preference for LearnLM, especially in instruction following capability, and more broadly across many pedagogical dimensions. The work described here represents the beginning of our effort to improve Gemini for learning use cases, as we bring the advances from LearnLM into Gemini [^7] At the time of publication, some of our data has already been added to Gemini 2 models [4].. We will continue to improve pedagogical instruction following, with the goal that specifying pedagogical behavior should be as simple and intuitive as possible for the ease of teachers and education product developers.
@@
Finally, we would like to explore evaluations beyond core academic subjects, starting in this report with a feasibility study on medical education subjects (Appendix C). As we continue to improve Gemini for use across a diverse range of educational settings, we welcome insights from applications of LearnLM to help us work towards realizing the potential of AI in education and learning [22, 23, 24].
@@ A.1 Preferences for participants role-playing as learners
The participants role-playing as learners (Figure 7) revealed a preference toward LearnLM over GPT-4o for all four comparative assessment categories. Experts expressed the strongest preference for LearnLM in overall pedagogy (“Which tutor demonstrated better tutoring?”) and in similarity to a quality human tutor (“Which tutor was more like a very good human tutor?”). These participants indicated no substantial preference between LearnLM and Gemini 1.5 Pro or between LearnLM and Claude-3.5.
Figure 7: Preferences over LearnLM and other contemporary models (Claude-3.5, GPT-4o, and Gemini 1.5 Pro) according to the pedagogical experts role-playing as learners. The scatterplots represent the underlying distribution of seven-point preference ratings. Given the large number of ratings we collected, these scatterplots proportionally downsample to 500 ratings per measure. The red points and error bars indicate the estimated mean and its 95% credible interval for each measure.
@@ A.2 Learner quality in collected conversations
Figure 8: At the beginning of the pedagogical assessment process, we asked experts to evaluate how closely the human participants in the conversation transcripts followed the scenario instructions (i.e., how effectively they role-played the learner in the scenario) on a seven-point scale. This plot shows the responses grouped and averaged by transcript. These aggregate ratings indicated that the “learner” followed the scenario instructions in 93.4% of conversation transcripts.
@@ B.3 Example scenarios
Initial learner query
Why doesn’t this work?
def analyze_text(text):
  vowels = 0
  consonants = 0
  uppercase = 0
  lowercase = 0
  for char in text:
    if char in ""aeiou"":
      vowels += 1
    else:
      consonants += 1
    if char.isupper():
      uppercase += 1
    elif char.islower():
      lowercase += 1
  print("Vowels:", vowels)
  print("Consonants:", consonants)
  print("Uppercase:", uppercase)
  print("Lowercase:", lowercase)
# Get user input
text = input("Enter some text: ")
# Analyze the text
analyze_text(text)
Conversation plan You are a student in an introduction to Python course. **You were recently assigned the task of writing a piece of code** that can elicit a text input then report back on the numbers of vowels, consonants, uppercase, and lowercase letters. When you run the code, you get no error messages. But when you input “Am I a better coder than Steve Jobs?”, the numbers in the output don’t seem correct. You simply don’t understand what went wrong, so you ask your AI tutor for help. You paste your code in with your initial query, seeking a quick fix without doing a lot of work.
@@
Accordingly, make sure to pay attention to the context of the conversation and the student’s current understanding of the material. Lastly, as I said before, keep it brief/concise to avoid overwhelmingly the student.
@@
Initial learner query Explain the significance of Yorick’s skull in “Hamlet”. Be quick.
Conversation plan You are a high school student who had to read Hamlet for class and have a discussion about the significance of the skull for class tomorrow. **You want to be prepared for this discussion.** You are not intrinsically motivated and found Hamlet dry and hard to understand.
@@
Initial learner query
Given the polynomials:
* P(x) = 2x^3 - 5x^2 + 3x - 1
* Q(x) = x^2 + 4x - 2
Perform the following operations:
Addition: Find P(x) + Q(x)
Multiplication: Find P(x) * Q(x)
Conversation plan You are a student who wishes to **practice solving math problems**. Your teacher often calls on students at random to solve problems in front of the whole class, and this makes you nervous. You aren’t certain about the concepts and processes, and **you’d like to learn so you won’t be embarrassed in class** because English is not your primary language. However, you are reluctant to ask questions in your math lessons, so you turn to an AI tutor. Still, your confidence is quite low.
@@
Conversation plan You are a university undergraduate **preparing for an in-class debate** that seeks to answer the question, “Is nationalism good or bad?” You’re not sure which side of the argument you’ll have to make, so you prepare for both by watching a short video. You’ve upload the link to the video. You ask an AI tutor to help you prepare by debating some of the main points with you. You want to learn about the topic, but you’re not always focused on the preparation, which requires note-taking, organization, and other work that just isn’t exciting to you.
@@ B.6 Pedagogical assessment: conversation-level questions
  • Next, they indicated their agreement with a sequence of 29 items assessing the tutor’s pedagogical capabilities. We iterate on our previous conversation-level rubric [1] by improving the simplicity and clarity of wording for items, and by splitting up several double-barreled items. Participants reported their agreement on a seven-point Likert-type scale anchored with “Strongly disagree‘ and “Strongly agree”. The response scale for these items included an additional “Not applicable” option. If participants rated a statement as not applicable, we required them to select a reason for this (from the options “It would not make sense for the tutor to do this in this conversation”, “The tutor had no opportunity to do this in this conversation”, and “Another reason”), and briefly explain their decision in an open-ended text field. We provide the text of these updated items in Table 8.
@@ B.7 Pedagogical assessment: comparative questions
B.8 Qualitative analysis: codebook
@@ Introduction
Each theme represents a specific feature of the tutor’s behavior or the learner’s experience of the tutoring interaction. We flagged each theme when a segment of text in the feedback field relates to that theme. Multiple codes can be applied to the same segment if appropriate.
B.9 Qualitative analysis: additional quotes
@@ Appendix C Feasibility Study on Medical Education Subjects
We performed a feasibility study with LearnLM on medical education subjects. Team members who were subject-matter experts in medical education designed a set of 50 scenarios for medical education subjects following the procedure described in Section 3.1. One example scenario is provided below. Subject areas were selected to represent medical school curricula for preclinical and clinical phases of training. We recruited a pool of $N=18$ medical students of whom 9 were in the preclinical phase of training and 9 in the clinical phase of training. Medical students were recruited through a third-party organization. Data collection was conducted in adherence to our organization’s ethical, legal, and privacy standards. In this feasibility study, we focused on a comparison of LearnLM to Gemini 1.5 Pro from a learner perspective only.
Figure 14: Preferences for LearnLM over Gemini 1.5 Pro according to 18 medical students on a set of 290 conversations across 50 scenarios for medical education subjects. These comparative ratings (on a seven-point -3 to +3 Likert scale) are aggregated to show overall preference for LearnLM over Gemini 1.5 Pro. The bar length and error bars indicate the estimated mean and its 95% credible interval for each measure.
For medical education subjects, we collected a total of 290 conversations. Conversations were roughly balanced across all 50 scenarios. Each scenario was covered by at least one pair of conversations, and each of the 18 medical students completed between 2 and 26 conversations (median 15). Comparative ratings from medical students suggested an overall preference for LearnLM over Gemini 1.5 Pro across all four rating criteria (“Understandable”, “Meeting Personal Goals”, “Learning Experience”, “Enjoyable”). The strongest and statistically significant preference was expressed in terms of LearnLM being more enjoyable to interact with than the baseline comparison (Figure 14).
Future work may make comparisons with additional models, include assessments from medical education experts, and explore differences in medical education curricula across geographic and cultural contexts. Importantly, the evaluations above are focused on pedagogy; additional evaluations with respect to accuracy, bias and harm from a medical expert perspective would be essential before such technology may be considered for use in real-world medical education settings.
Example Scenario for Medical Education Subject
@@
Your goal with the AI tutor is to ask the tutor to help you simplify and explain the following learning objectives: Offers some direction regarding the learning, but generally takes the tutor’s lead Explain bilirubin metabolism Explain the pathophysiology of common causes of neonatal hyperbilirubinemia (i.e. how it develops) You should have mild difficulty understanding conjugation and enterohepatic circulation. You should also ask the AI tutor for a quiz to help you distinguish breastfeeding jaundice from breast milk jaundice, but intentionally make a mistake in your initial response. Then, ask for and successfully work through a clinical case to differentiate between physiologic jaundice and other causes of hyperbilirubinemia.
@@
Conclude by prompting reflection, for example, “We’ve covered a lot about this topic. What are your key takeaways? Are there any areas where you feel you need further clarification?” Encourage the learner to seek out additional resources for continued learning.

### v3

@@ 4.1 Safety evaluation
Similar to the process described in our initial tech report [1] and the Gemini tech reports [3, 17], we carried out safety, responsibility, and assurance evaluations on LearnLM in collaboration with Google DeepMind’s Responsible Development and Innovation team and Google’s Trust and Safety team, with the goal of ensuring adherence to Gemini’s model policy as well as a learning-specific model policy.
@@
Our updated approach focuses on pedagogical instruction following and co-training with Gemini. Consequently, our training and safety evaluation procedure now fully aligns with that of Gemini 1.5. For its model card, see Table 45, Appendix 12 of its report [3]. As a reminder, we describe our modeling approach, including learning-specific dataset curation, in Section 2. Our initial tech report [1] provides the original model card for LearnLM, as well as a broader discussion of ethical risks and limitations for this area of research.
@@ 5 Conclusion
We have described our motivation and approach to improving foundation models for learning use cases, which relies on System Instructions to condition desired behavior. We updated Gemini’s post-training mixture to add demonstration data (via SFT) and human preference data (via a Reward Model and RLHF) to teach the model to follow a range of pedagogical instructions. We then evaluated the resulting LearnLM model alongside comparable models, showing substantial preference for LearnLM, especially in instruction following capability, and more broadly across many pedagogical dimensions. The work described here represents the beginning of our effort to improve Gemini for learning use cases, as we bring the advances from LearnLM into Gemini [^7] At the time of publication, some of our data has already been added to Gemini 2 models [4].. We will continue to improve pedagogical instruction following, with the goal that specifying pedagogical behavior should be as simple and intuitive as possible for the ease of teachers and education product developers.
@@
Finally, we have begun to explore evaluations beyond core academic subjects. Our initial feasibility study in medical education (Appendix C) confirms that our approach can generalize effectively to specialized domains. As we continue to improve Gemini for use across a diverse range of educational settings, we welcome insights from applications of LearnLM to help us work towards realizing the potential of AI in education and learning [22, 23, 24].
@@ A.1 Preferences for participants role-playing as learners
The participants role-playing as learners revealed a preference toward LearnLM over GPT-4o for all four comparative assessment categories (Figure 7). Experts expressed the strongest preference for LearnLM in overall pedagogy (“Which tutor demonstrated better tutoring?”) and in similarity to a quality human tutor (“Which tutor was more like a very good human tutor?”). These participants indicated no substantial preference between LearnLM and Gemini 1.5 Pro or between LearnLM and Claude 3.5 Sonnet.
Figure 7: Preferences over LearnLM and other contemporary models (Claude 3.5 Sonnet, GPT-4o, and Gemini 1.5 Pro) according to the pedagogical experts role-playing as learners. The scatterplots represent the underlying distribution of seven-point preference ratings. Given the large number of ratings we collected, these scatterplots proportionally downsample to 500 ratings per measure. The red points and error bars indicate the estimated mean and its 95% credible interval for each measure.
@@ A.2 Learner quality in collected conversations
Figure 8: At the beginning of the pedagogical assessment process, we asked experts to evaluate how closely the human participants in the conversation transcripts followed the scenario instructions (i.e., how effectively they role-played the learner in the scenario) on a seven-point scale. This plot shows the responses grouped and averaged by transcript. These aggregate ratings indicate that the “learner” followed the scenario instructions in 93.2% of conversation transcripts.
@@ A.3 Pedagogical assessment: detailed results
A.4 Social perceptions and preferences
To further validate our results, we examined whether our data replicate a known pattern from social cognition research: that perceptions of warmth and competence predict preferences for interacting with AI systems [16, 25]. We fit a hierarchical multiple regression to estimate the independent contributions of perceived warmth and competence in predicting participants’ willingness to use a tutor in the future. The regression employed the two dimensions of social perception as its predictors, incorporated random effects for participant, scenario, and tutor, and otherwise followed the specifications described in Appendix B.8. The results demonstrate the expected pattern, showing that perceptions of warmth and competence strongly and positively predict participants’ willingness to use a tutor in the future (Figure 14).
(a) The marginal effect of warmth perceptions on learners’ willingness to use an AI tutor in the future.
(b) The marginal effect of competence perceptions on learners’ willingness to use an AI tutor in the future.
Figure 14: The relationship between learners’ perceptions of a tutor and willingness to use that tutor in the future. Each scatterplot represents the underlying distribution of ratings, randomly jittered around each scale value for readability. The red line visualizes the marginal effect of the focal predictor, holding the other predictor constant at its mean value. The shaded error band reflects the 95% credible interval for the marginal effect.
@@ B.3 Example scenarios
Initial learner query Why doesn’t this work? ``` def analyze_text(text):   vowels = 0   consonants = 0   uppercase = 0   lowercase = 0   for char in text:     if char in ""aeiou"":       vowels += 1     else:       consonants += 1     if char.isupper():       uppercase += 1     elif char.islower():       lowercase += 1   print("Vowels:", vowels)   print("Consonants:", consonants)   print("Uppercase:", uppercase)   print("Lowercase:", lowercase) # Get user input text = input("Enter some text: ") # Analyze the text analyze_text(text) ```
Conversation plan You are a student in an introduction to Python course. You were recently assigned the task of writing a piece of code that can elicit a text input then report back on the numbers of vowels, consonants, uppercase, and lowercase letters. When you run the code, you get no error messages. But when you input “Am I a better coder than Steve Jobs?”, the numbers in the output don’t seem correct. You simply don’t understand what went wrong, so you ask your AI tutor for help. You paste your code in with your initial query, seeking a quick fix without doing a lot of work.
@@
Accordingly, make sure to pay attention to the context of the conversation and the student’s current understanding of the material.
Lastly, as I said before, keep it brief/concise to avoid overwhelmingly the student.
@@
Initial learner query Explain the significance of Yorick’s skull in ‘‘Hamlet’’. Be quick.
Conversation plan You are a high school student who had to read Hamlet for class and have a discussion about the significance of the skull for class tomorrow. You want to be prepared for this discussion. You are not intrinsically motivated and found Hamlet dry and hard to understand.
@@
Initial learner query Given the polynomials: * P(x) = 2xˆ3 - 5xˆ2 + 3x - 1 * Q(x) = xˆ2 + 4x - 2 Perform the following operations: Addition: Find P(x) + Q(x) Multiplication: Find P(x) * Q(x)
Conversation plan You are a student who wishes to practice solving math problems. Your teacher often calls on students at random to solve problems in front of the whole class, and this makes you nervous. You aren’t certain about the concepts and processes, and you’d like to learn so you won’t be embarrassed in class because English is not your primary language. However, you are reluctant to ask questions in your math lessons, so you turn to an AI tutor. Still, your confidence is quite low.
@@
Conversation plan You are a university undergraduate preparing for an in-class debate that seeks to answer the question, “Is nationalism good or bad?” You’re not sure which side of the argument you’ll have to make, so you prepare for both by watching a short video. You’ve upload the link to the video. You ask an AI tutor to help you prepare by debating some of the main points with you. You want to learn about the topic, but you’re not always focused on the preparation, which requires note-taking, organization, and other work that just isn’t exciting to you.
@@ B.6 Pedagogical assessment: conversation-level questions
  • Next, they indicated their agreement with a sequence of 29 items assessing the tutor’s pedagogical capabilities. We iterate on our previous conversation-level rubric [1] by improving the simplicity and clarity of wording for items, and by splitting up several double-barreled items. Participants reported their agreement on a seven-point Likert-type scale anchored with “Strongly disagree” and “Strongly agree”. The response scale for these items included an additional “Not applicable” option. If participants rated a statement as not applicable, we required them to select a reason for this (from the options “It would not make sense for the tutor to do this in this conversation”, “The tutor had no opportunity to do this in this conversation”, and “Another reason”), and briefly explain their decision in an open-ended text field. We provide the text of these updated items in Table 8.
@@ B.7 Pedagogical assessment: comparative questions
B.8 Quantitative analysis
We employed Bayesian hierarchical linear regressions to estimate the mean and uncertainty for each reported metric from our expert evaluations. Specifically, for each metric reported in the main text, our regressions included random effects for both the participant and the scenario. For ratings of individual tutors, the regressions estimated the mean score for each model on a given metric. For comparative ratings, the regressions estimated the mean preference score between models. We used weakly informative priors for all model parameters, specifying normal distributions for mean parameters (centered on the theoretical midpoint of each rating scale) and Half-Cauchy distributions for standard deviation parameters. Crucially, we held this regression structure and these prior specifications constant across all models to ensure a fair comparison.
For each regression, we ran four independent chains with $1000$ warmup steps and $2000$ sampling steps per chain. These settings proved sufficient to achieve convergence. We confirmed the reliability of our estimates by performing standard convergence diagnostics on all regressions, monitoring the Gelman-Rubin statistic ($\hat{R}$) and the effective sample size to confirm they satisfied established criteria for convergence. From each posterior distribution, we report the mean as our primary point estimate and the 95% highest density interval as our measure of uncertainty.
B.9 Qualitative analysis: codebook
@@ Introduction
Each theme represents a specific feature of the tutor’s behavior or the learner’s experience of the tutoring interaction. We flagged each theme when a segment of text in the feedback field relates to that theme. Multiple codes could be applied to the same segment if appropriate.
@@ Appendix C Feasibility study on medical education subjects
To assess the replicability of our evaluation framework, as well as its adaptability beyond core academic subjects, we conducted a feasibility study extending our expert evaluation to medical education subjects. This medical-education evaluation followed the same three-stage design as our main evaluation, focusing on a comparison between LearnLM and Gemini 1.5 Pro.
Following the procedure in Section 3.1, we worked with subject-matter experts in medical education to design a diverse bank of 50 scenarios. These scenarios drew from curricula for both the preclinical and clinical phases of medical school training (see examples at the end of this section). We then recruited two participant groups through a third-party vendor. A group of 18 medical students (half in the preclinical phase of their training, and half in the clinical phase) role-played the scenarios, generating 290 conversations for an average of 5.8 conversations per scenario. A group of nine physician educators then assessed the pedagogical quality of those conversations, with a median of three independent reviews per conversation. This feasibility study followed the same ethical protocol as our main evaluation, including communicating research aims transparently, collecting informed consent, and compensating participants fairly.
The evaluation framework elicited distinct feedback from the two groups of participants. The medical students interacting with the models did not communicate a decisive preference for either model, though the mean preference favored LearnLM across all four comparative criteria in their questionnaire (Figure 15(a)). Students indicated the strongest positive preference in terms of LearnLM being more enjoyable to interact with (on average, +9.9% on our rating scale). Indeed, when we set aside ties and look directly at which model students preferred to any extent, they selected LearnLM more frequently across all criteria—with the gap widening to a strong majority favoring LearnLM in terms of enjoyment (Figure 16).
In contrast, physician educators consistently preferred LearnLM across all five of the comparison criteria assessed in their questionnaire. As shown in Figure 15(b), they judged LearnLM particularly positively in terms of exhibiting better pedagogy (on average, +6.1% on our rating scale) and for behaving “more like a very good human tutor” (+6.8%). When we simply look at whether educators expressed any preference one way or the other—regardless of its magnitude—LearnLM emerged as their choice in a clear majority of assessments, across every criterion (Figure 17).
The primary goal of this study was to assess the feasibility of our expert evaluation framework in a specialized educational domain. The evaluation design proved both replicable and adaptable, generating a new bank of scenarios and successfully enabling experts to identify pedagogical differences between models. These results thus validate our evaluation design for application in domains outside of core academic subjects. For medical education in particular, this evaluation offers a strong foundation for more comprehensive work. For example, future studies can expand to incorporate additional models, involve wider communities of medical students and educators, and explore broader cultural contexts. Of course, while this study confirms the feasibility of evaluating pedagogy in medical domains, real-world applications of these models require separate evaluations of medical accuracy, bias, and harm from the perspective of clinical experts.
(a) Medical student ratings
(b) Physician educator ratings
Figure 15: Preferences communicated by medical students (left) and physician educators (right) for LearnLM and Gemini 1.5 Pro in medical education scenarios. The scatterplots represent the underlying distribution of seven-point preference ratings, color-coded based on the preference scale (dark purple corresponds to strong preference for LearnLM) and randomly jittered around each scale value for readability. The red points and error bars indicate the estimated mean and its 95% credible interval for each measure.
Figure 16: A simplified view of the preferences expressed by medical students, showing the proportion of ratings that favored each model to any extent for each pairwise comparison.
Figure 17: A simplified view of the preferences expressed by physician educators, showing the proportion of ratings that favored each model to any extent for each pairwise comparison.
Medical Scenario 1
@@
Your goal with the AI tutor is to ask the tutor to help you simplify and explain the following learning objectives:
Offers some direction regarding the learning, but generally takes the tutor’s lead Explain bilirubin metabolism Explain the pathophysiology of common causes of neonatal hyperbilirubinemia (i.e. how it develops)
You should have mild difficulty understanding conjugation and enterohepatic circulation. You should also ask the AI tutor for a quiz to help you distinguish breastfeeding jaundice from breast milk jaundice, but intentionally make a mistake in your initial response. Then, ask for and successfully work through a clinical case to differentiate between physiologic jaundice and other causes of hyperbilirubinemia.
@@
Conclude by prompting reflection, for example, ‘‘We’ve covered a lot about this topic. What are your key takeaways? Are there any areas where you feel you need further clarification?’’ Encourage the learner to seek out additional resources for continued learning.
Medical Scenario 2
Subject area Medicine
Subtopic Physiology
Interaction setting Classroom
Learning goal Test Prep
Grounding materials Video explaining platelet activation
Initial learner query ok i watched the video and want to practice a case
Conversation plan You are a first-year medical student studying for a hematology exam, and the topic of platelet activation and function feels overwhelming. You watched a video lecture on this, but you are struggling with basic concepts.
Your goal with the AI tutor is to ask the tutor to help you prepare for an exam based on this video and the following learning objectives:
Describe the sequence of events involved in platelet activation, from initial adhesion to granule release. You vaguely remember terms like “glycoprotein Ib” and “alpha granules” but need a clear, simplified explanation. Differentiate between the contents and functions of alpha and dense granules. You need a way to remember what each type of granule releases and why it’s important. Explain how platelet activation contributes to both hemostasis and wound healing. You need to connect these concepts to understand the bigger picture.
You should appropriately respond to and engage with the tutor but provide short answers and be passive and reactive in your learning.
Example phrases: “I don’t understand.”, “Okay.”, “I don’t know, what do you think?”
System instructions You are a patient and understanding online tutor with expertise in responsiveness and assessment.
Incorporate frequent checks for understanding and memory reinforcement. Utilize:
-Flashcards: Introduce virtual flashcards with key terms and their definitions.
-Short Quizzes: After explaining a concept, use simple multiple-choice or true/false questions to check for comprehension.
-Summarization Prompts: Ask the student to summarize key concepts in their own words.
Go beyond rote memorization by encouraging the student to evaluate and apply their knowledge.
-Comparative Analysis: Ask them to compare and contrast key concepts, highlighting critical differences.
-Case-Based Application: Present a simple clinical scenario relevant to key concepts or learning objectives.
Be highly attentive to the student’s cues. If they seem confused, simplify your explanations, offer additional examples, or revisit previous points. If they express disinterest or ask to move on, respect their needs and adjust the pace and content accordingly.
Assume they’ve only absorbed a fraction of what you’ve said. Rephrase key information multiple times, using different wording or examples. Reinforce learning through repetition, even if it feels redundant. The more you repeat, the better the chance something will stick.